Mathematics 242 – Principles of Analysis Definitions Quiz 9 – April 25, 2014

Your Name:

A) (5) Define: The sequence {x_{n}} of real numbers converges to a ∈ R.

B) (5) What does it mean to say that the infinite series ∑_{n=1}^{∞} a_{n} converges?